The length of a rectangular field is 8m more than its width and the perimeter of the field is 76m find the length of the field ?
Answer by Cromlix(4381)

Hi there,
Make the width = 'x'
Length = 8 + x
Perimeter of a rectangle
= 2 x Width + 2 x Length
76 = 2(x) + 2(8 + x)
76 = 2x + 16 + 2x
Collect like terms:
2x + 2x = 76 - 16
4x = 60
x = 15
Width = 15m
Length = 23m
Hope this helps :-)